# A phase-space view of vibrational energies without the Born-Oppenheimer framework

### v1

@@ Abstract
I Introduction
I.1 The Hydrogen Atom
The Born-Oppenheimer (BO) framework[1, 2, 3] is the standard approach to understanding chemistry, upon which the vast majority of chemical intuition is built. The basic ansatz of BO theory is that, because nuclei are slow and electrons are fast, one can separate the total Hamiltonian into two terms, the nuclear kinetic energy and the electronic Hamiltonian:
@@ I.1 The Hydrogen Atom
One critical limitation of BO theory can be identified by considering the hydrogen atom, the most well understood problem in quantum mechanics. The Hamiltonian is simply the kinetic energy of the electron plus the kinetic energy of the proton plus their interaction:
$\hat{H}_{\rm Hydrogen}=\frac{\hat{\bm{P}}_{\rm n}^{2}}{2M}+\frac{\hat{\bm{p}}_{\rm e}^{2}}{2m}-\frac{1}{|{\bm{\hat{r}}_{\rm e}-\bm{\hat{R}}_{\rm n}}|}.$ (2)
@@
As can be found in any textbook[4], the way to solve this Hamiltonian is to change variables from $\hat{\bm{R}}_{\rm n}$ and $\hat{\bm{r}}_{\rm e}$ to the relative position and center of mass coordinates:
$\hat{\bm{r}}=\hat{\bm{r}}_{\rm e}-\hat{\bm{R}}_{\rm n},\quad\hat{\bm{R}}=\frac{M\hat{\bm{R}}+m_{\rm e}\hat{\bm{r}}}{M+m_{\rm e}}.$ (3)
In so doing, the Hamiltonian becomes:
$\hat{H}_{\rm com}=\frac{\hat{\bm{P}}^{2}}{2(m_{\rm e}+M)}+\frac{\hat{\bm{p}}^{2}}{2\mu}-\frac{1}{|\hat{\bm{r}}|}$ (4)
@@
with the reduced mass $\mu=(M^{-1}+m_{e}^{-1})^{-1}$, and the stationary energies for the system are the center of mass kinetic energy plus the quantized bound state energy of the atom:
$\epsilon_{n}^{\rm exact}(\bm{P})=\frac{\bm{P}^{2}}{2(M+m_{\rm e})}-\frac{\mu e^{4}}{8\epsilon_{0}^{2}h^{2}}\frac{1}{n^{2}}.$ (5)
@@
Note that, within the BO (clamped nuclei) approximation, the result is slightly different if one diagonalizes $\hat{H}_{el}$:
$\epsilon_{n}^{\rm BO}(\bm{P})=\frac{\bm{P}^{2}}{2M}-\frac{m_{\rm e}e^{4}}{8\epsilon_{0}^{2}h^{2}}\frac{1}{n^{2}}.$ (6)
@@
Now, on the one hand, one might argue that the energy expressions are not that different between Eq. 5 and Eq. 6. That being said, if we ignore energy for a moment, there is a conceptual problem insofar as the fact that, unless one changes coordinates, the electronic momentum of the ground state must be zero[5, 6, 7]; in other words, even if the hydrogen atoms is moving, the electronic momentum is still calculated to be zero (which is incorrect). Moreover, in the context of energy, if we were to consider a system of $N$ isolated hydrogen atoms, the total difference in energy becomes $N$ times larger. Furthermore, if one were to bring those $N$ atoms closer together, so that they could interact and form hydrogen molecules or solids, there is no reason to assume that the problem of freezing nuclei (rather than choosing optimal coordinates) would go away[8, 9, 10, 11, 12, 13, 14] – in fact, one might expect that with more and more interacting nuclei, the problems of BO theory would only become more daunting and more difficult to solve.
@@
Within the context of molecular quantum chemistry, the problem above is often addressed in the literature by separating out the center of mass, in the spirit of Eq. 4 above[15, 16, 17]. Mathematically, one can formally change coordinates and separate out the total center of mass by choosing Jacobi coordinates for the nuclei and relating all electrons to the nuclear center of mass:
$\hat{\bm{R}}_{\rm n,com}=\frac{\sum_{A}M_{A}{\hat{\bm{R}}_{A}}}{\sum_{A}M_{A}},\quad\hat{\bm{r}}_{i}=\hat{\bm{r}}_{i}-\hat{\bm{R}}_{\rm n,com}.$ (7)
@@
It is crucial to emphasize that these transformations do not solve the problem above – especially in the context of condensed phase chemistry with many molecules. In particular, even for a pair of H atoms separated far from each other, e.g. one on the earth and one on the moon, the transformation in Eq. 7 introduces one coordinate (the nuclear center of mass) which is somewhere in the exosphere, but does not succeed at recovering the correct ground state energy.
I.2 Review of Electronic Phase-Space Hamiltonians
@@
New approaches that go beyond BO theory are clearly needed for accurate quantum chemistry[18, 19, 20]. With this in mind, we have recently proposed a phase-space approach to treat the exact problem above. To motivate this approach, consider the Hamiltonian in Eq. 2 as rewritten in an atomic adiabatic basis (i.e. after diagonalization of $\hat{U}\hat{H}_{\rm el}\hat{U}^{\dagger}=\hat{V}_{\rm ad}$):
@@
where $\hat{\bm{p}}_{\rm e}^{\rm ad}=\hat{U}\hat{\bm{p}}_{e}\hat{U}^{\dagger}$. If one now makes the ansatz that the nuclei are classical, one can construct a phase-space electronic Hamiltonian which is a function of the parameters (not operators) $\bm{R}_{\rm n}$ and $\bm{P}_{\rm n}$ and transform it back to the original basis:
$\hat{H}_{\rm PS}=\frac{{\bm{P}}_{\rm n}^{2}}{2M}+\hat{H}_{\rm el}-\frac{{\bm{P}}_{\rm n}\cdot\hat{\bm{p}}_{\rm e}}{M}+\frac{\hat{\bm{p}}_{\rm e}^{2}}{2M}.$ (9)
A simple analysis shows that $\hat{H}_{\rm PS}$ represents a boost of the electron with nuclear momentum ${\bm{P}}_{\rm n}$, and its corresponding eigenvalues are identical to the exact energies for a hydrogen atom:
$\epsilon_{n}^{\rm PS}({\bm{P}}_{\rm n})=\frac{\bm{P}_{\rm n}^{2}}{2M}-\frac{\mu e^{4}}{8\epsilon_{0}^{2}h^{2}}\frac{1}{n^{2}}$ (10)
(provided we equate ${\bm{P}}_{\rm n}$ [which captures the nuclear momentum] with $\bm{P}$ [which is formally the center of mass momentum]).
This finding is very encouraging because it offers an approach for a system with $N$ isolated hydrogen atoms. In such a case, all we must do is to partition the electronic momentum into different operators, one around each atom, so that one can effectively perform $N$ different local changes of coordinate (around each atom):
@@
where $\hat{\theta}_{A}(\hat{\bm{r}})$ is a partition of space:
@@
Note that while there are many ways to choose the space partition function $\hat{\theta}_{A}$, a natural option is to choose smooth functions that are localized on atoms (e.g., a combination of Gaussian functions, see supporting information [SI]).
One can then write:
@@
This is almost the final form of $\hat{H}_{\rm PS}$, but there are two more important points. First, one can show that if one diagonalizes Eq. 14, then classical dynamics propagated along the resulting energy surfaces will automatically conserve linear momentum. If one wishes to conserve angular momentum, one must also include another term to account for angular momentum[21]:
A preliminary analysis of the relevant equations of motion is given in the SI.
Second, the form of the phase-space Hamiltonian in Eq. 14 is reasonable insofar as, if one diagonalizes $\hat{H}_{\rm el}(\bm{R})$, the exact Hamiltonian in the basis of electronic eigenstates is of the form:
@@
where $\hat{\bm{d}}_{A}$ is the derivative coupling and $\hat{V}_{\rm ad}$ is the diagonal matrix of eigenvalues. Thus, one concludes that the phase-space Hamiltonian is the natural electronic object that emerges when we make the nuclear position and momentum classical,
$\hat{\bm{R}}\rightarrow\bm{R},\quad\hat{\bm{P}}\rightarrow\bm{P}$ (18)
and one further approximates the derivative coupling by an operator $\hat{\bm{d}}_{A}\rightarrow\hat{\bm{\Gamma}}_{A}=\hat{\bm{\Gamma}}^{\prime}_{A}+\hat{\bm{\Gamma}}^{\prime\prime}_{A}$. By construction, $\hat{\bm{\Gamma}}_{A}$ satisfies[21, 22, 23]

### v2

@@ Abstract
1 Introduction
1.1 The Hydrogen Atom
The Born-Oppenheimer (BO) framework1, 2, 3 is the standard approach to understanding chemistry, upon which the vast majority of chemical intuition is built. The basic ansatz of BO theory is that, because nuclei are slow and electrons are fast, one can separate the total Hamiltonian into two terms, the nuclear kinetic energy and the electronic Hamiltonian:
@@ 1.1 The Hydrogen Atom
One critical limitation of BO theory can be identified by considering the hydrogen atom, the most well-understood problem in quantum mechanics. The Hamiltonian is simply the kinetic energy of the electron plus the kinetic energy of the proton plus their interaction:
$\hat{H}_{\rm Hydrogen}$ $=$ $\frac{\hat{\bm{P}}_{\rm n}^{2}}{2M}+\hat{H}^{\rm el}_{\rm Hydrogen}$ (2)
$\hat{H}^{\rm el}_{\rm Hydrogen}$ $=$ $\frac{\hat{\bm{p}}_{\rm e}^{2}}{2m_{\rm e}}-\frac{e^{2}}{4\pi\epsilon_{0}|{\bm{\hat{r}}_{\rm e}-\bm{\hat{R}}_{\rm n}}|}.$ (3)
1.2 Standard Approach: Separate out the center of mass
@@
As can be found in any textbook4, the way to solve this Hamiltonian is to change variables from $\hat{\bm{R}}_{\rm n}$ and $\hat{\bm{r}}_{\rm e}$ to the relative position and center of mass coordinates:
$\hat{\bm{r}}=\hat{\bm{r}}_{\rm e}-\hat{\bm{R}}_{\rm n},\quad\hat{\bm{R}}=\frac{M\hat{\bm{R}}_{\rm n}+m_{\rm e}\hat{\bm{r}}}{M+m_{\rm e}}.$ (4)
with the corresponding momentum operators $\hat{\bm{p}}$ and $\hat{\bm{P}}$ in the center of mass coordinates as well:
$\hat{\bm{p}}=\hat{\bm{P}_{n}}-\hat{\bm{p}_{e}},\quad\hat{\bm{P}}=\hat{\bm{P}_{n}}+\hat{\bm{p}_{e}}.$ (5)
In doing so, the Hamiltonian becomes:
$\hat{H}_{\rm com}=\frac{\hat{\bm{P}}^{2}}{2(M+m_{\rm e})}+\frac{\hat{\bm{p}}^{2}}{2\mu}-\frac{e^{2}}{4\pi\epsilon_{0}|\hat{\bm{r}}|}$ (6)
@@
Here, the reduced mass is $\mu=(M^{-1}+m_{e}^{-1})^{-1}$. The stationary energies for the system are then the center of mass kinetic energy plus the quantized bound state energy of the atom:
$E_{n}^{\rm exact}(\bm{P})=\frac{\bm{P}^{2}}{2(M+m_{\rm e})}-\frac{\mu e^{4}}{8\epsilon_{0}^{2}h^{2}}\frac{1}{n^{2}}.$ (7)
@@
Note that, within the BO (clamped-nuclei) approximation, the result is slightly different if one diagonalizes $\hat{H}_{\rm el}$:
$E_{n}^{\rm BO}(\bm{P}_{\rm n})=\frac{\bm{P}_{\rm n}^{2}}{2M}-\frac{m_{\rm e}e^{4}}{8\epsilon_{0}^{2}h^{2}}\frac{1}{n^{2}}.$ (8)
@@
Now, one might argue that the energy expressions are not that different between Eq. 7 and Eq. 8. That being said, if we ignore the raw energy difference for a moment, there is a conceptual problem insofar as the fact that, unless one changes coordinates, the electronic momentum of the ground state must be zero5, 6, 7; in other words, even if the hydrogen atom is moving, the electronic momentum is still calculated to be zero (which is incorrect). Moreover, in the context of energy, if we were to consider a system of $N$ isolated hydrogen atoms, the total difference in energy becomes $N$ times larger. Furthermore, if one were to bring those $N$ atoms closer together, so that they could interact and form hydrogen molecules or solids, there is no reason to assume that the problem of freezing nuclei (rather than choosing optimal coordinates) would go away8, 9, 10, 11, 12, 13 – in fact, one might expect that with more and more interacting nuclei, the problems of BO theory would only become more daunting and more difficult to solve.
@@
Within the context of molecular quantum chemistry, it might appear that the problem above can be addressed by separating out the center of mass, in the spirit of Eq. 6 above14, 15, 16. Mathematically, for a system with an arbitrary number of electrons and nuclei, one can formally change coordinates and separate out the total center of mass by choosing Jacobi coordinates for the nuclei $\{\bm{R}_{i}\}$ and defining new electrons coordinates $\bm{r}_{a}^{\prime}$ relative to the nuclear center of mass ${\bm{R}}_{\rm com}$:
$\hat{\bm{R}}_{\rm com}=\frac{\sum_{A}M_{A}{\hat{\bm{R}}_{A}}}{\sum_{A}M_{A}},\quad\hat{\bm{r}}_{a}^{\prime}=\hat{\bm{r}}_{a}-\hat{\bm{R}}_{\rm com}.$ (9)
Note that here and throughout the paper, we will use Latin letters $A,B,C$ to label different atoms, $a,b,c$ to label different electrons and Greek letters $\alpha,\beta,\gamma$ for Cartesian coordinates. By defining a coordinate transformation $(\bm{R}_{A},\bm{r}_{a})\to({\bm{R}}_{\rm com},\bm{R}_{i},\bm{r}_{a}^{\prime})$, the molecular Hamiltonian in Eq. 1 becomes:
$\hat{H}=\frac{\hat{\bm{P}}^{2}_{\rm com}}{2M_{\rm tot}}+\sum_{i}\frac{\hat{\bm{P}}_{i}^{2}}{2\mu_{i}}+\sum_{a}\frac{\hat{\bm{p}^{\prime}}^{2}_{a}}{2m_{\bm{e}}}+\sum_{a,b}\frac{\hat{\bm{p}^{\prime}}_{a}\cdot\hat{\bm{p}^{\prime}}_{b}}{2M_{\rm n}}$ (10)
with the total nuclear mass $M_{\rm n}=\sum_{A}{M_{A}}$, the total molecular mass $M_{\rm tot}=M_{\rm n}+\sum_{a}{m_{\rm e}}$ and the reduced mass $\mu_{i}$ corresponding to the $i$-th internal nuclear Jacobi coordinate $\bm{R}_{i}$. The last term in Eq. 10 is often referred to as the mass polarization term 17 and a detailed definition of the internal coordinate $\bm{R}_{i}$ can be found in any standard textbooks18.
@@
Now, it is crucial to emphasize that the transformations leading to Eq. 10 above do not solve the problem of electronic momentum – especially in the context of condensed phase chemistry with many molecules. In particular, consider a pair of H atoms separated far from each other, e.g. one on the earth and one on the moon. The correct eigenspectrum is characterized by two good quantum numbers (one for each atomic level) plus the kinetic energy of each hydrogen atom:
$E_{n_{1},n_{2}}^{\rm exact}(\bm{P}_{1},\bm{P}_{2})=\frac{\bm{P}_{1}^{2}}{2(M+m_{\rm e})}+\frac{\bm{P}_{2}^{2}}{2(M+m_{\rm e})}-\frac{\mu e^{4}}{8\epsilon_{0}^{2}h^{2}}\frac{1}{n_{1}^{2}}-\frac{\mu e^{4}}{8\epsilon_{0}^{2}h^{2}}\frac{1}{n_{2}^{2}}.$ (11)
Note that the transformation in Eq. 9 does not succeed at recovering this (correct) electronic eigen-spectrum, but instead introduces an irrelevant coordinate (the nuclear center of mass) which is somewhere in the exosphere. As a side note, in this article, we will ignore all relativistic and quantum electrodynamics effects and focus on effects beyond BO theory due to the finite nuclear mass.
2 Phase-Space Electronic Hamiltonian Theory
2.1 Phase-Space Interpretation of the Hydrogen Atom
@@
New approaches that go beyond BO theory are clearly needed for accurate quantum chemistry19, 20. With this in mind, we have recently proposed a phase-space approach to treat the exact problem above. To motivate this approach, consider the hydrogen Hamiltonian in Eq. 2. Let $\hat{U}$ be the unitary operator that diagonalizes the electronic Hamiltonian in Eq. 3:
$\hat{U}\hat{H}^{\rm el}_{\rm Hydrogen}\hat{U}^{\dagger}$ $=$ $\hat{V}_{\rm ad}$ (12)
${\hat{H}}_{\rm Hydrogen}^{\rm el}\ket{\psi_{n}}$ $=$ $E_{n}^{\rm BO}\ket{\psi_{n}}$ (13)
As is well known21, the total Hamiltonian in this basis
$\hat{H}_{\rm ad}=\hat{U}\hat{H}_{\rm hydrogen}\hat{U}^{\dagger}$ (14)
has matrix elements that are given by:
$H_{{\rm ad},mn}=\sum_{k}\frac{(\hat{\bm{P}}_{\rm n}\delta_{mk}-i\hbar{\bm{d}}_{mk})(\hat{\bm{P}}_{\rm n}\delta_{kn}-i\hbar{\bm{d}}_{kn})}{2M}+E_{n}^{\rm BO}\delta_{mn}$ (15)
where the non-adiabatic coupling vector ${\bm{d}_{mk}}=\bra{\psi_{m}}\bm{\nabla}_{\bm{R}_{\rm n}}\ket{\psi_{k}}$.
Now, because of the translational invariance of the adiabatic basis {|$\{\ket{\psi_{n}}\}$, i.e.
$\frac{\hbar}{i}\left(\frac{\partial}{\partial{\bm{r}}_{e}}+\frac{\partial}{\partial{\bm{R}}_{n}}\right)\ket{\psi_{n}}=\left(\hat{\bm{P}}_{\rm n}+\hat{\bm{p}}_{\rm e}\right)\ket{\psi_{n}}=0,$ (16)
it follows that we can express Eq. 15 as:
$H_{{\rm ad},mn}=\sum_{k}\frac{(\hat{\bm{P}}_{\rm n}\delta_{mk}-{\bm{p}}_{{\rm e},mk})(\hat{\bm{P}}_{\rm n}\delta_{kn}-{\bm{p}}_{{\rm e},kn})}{2M}+E_{n}^{\rm BO}\delta_{mn}$ (17)
with ${\bm{p}}_{{\rm e},mk}=\bra{\psi_{m}}\hat{\bm{p}}_{\rm e}\ket{\psi_{k}}.$ Thus, we can write this operator in basis-free form as:
@@
where $\hat{\bm{p}}_{\rm e}^{\rm ad}=\hat{U}\hat{\bm{p}}_{e}\hat{U}^{\dagger}$. Up to this point, there has been no approximation: Eq. 18 is an exact representation of the hydrogen atom Hamiltonian that sets the stage for a phase-space approach.
If one now makes the ansatz that the nuclei are classical, one can construct a phase-space electronic Hamiltonian which is a function of the parameters (not operators) $\bm{R}_{\rm n}$ and $\bm{P}_{\rm n}$:
$\hat{H}^{\rm ad}_{\rm PS}({\bm{R}}_{\rm n},{\bm{P}}_{\rm{}_{n}})=\frac{{\bm{P}}_{\rm n}^{2}}{2M}+\hat{V}_{\rm ad}({\bm{R}}_{\rm n})-\frac{{\bm{P}}_{\rm n}\cdot\hat{\bm{p}}_{\rm e}^{\rm ad}+\hat{\bm{p}}_{\rm e}^{\rm ad}\cdot{\bm{P}}_{\rm n}}{2M}+\frac{(\hat{\bm{p}}_{\rm e}^{\rm ad})^{2}}{2M}.$ (19)
At this point, and only for practical purposes, the most helpful next step is to transform the Hamiltonian back to the original (presumably diabatic) basis:
$\hat{H}_{\rm PS}$ $=$ $\hat{U}^{\dagger}\hat{H}^{\rm ad}_{\rm PS}\hat{U}$ (20)
$=$ $\frac{{\bm{P}}_{\rm n}^{2}}{2M}+\hat{H}_{\rm el}-\frac{{\bm{P}}_{\rm n}\cdot\hat{\bm{p}}_{\rm e}}{M}+\frac{\hat{\bm{p}}_{\rm e}^{2}}{2M}.$ (21)
A simple analysis shows that $\hat{H}_{\rm PS}$ represents a boost of the electron with nuclear momentum ${\bm{P}}_{\rm n}$,
$\hat{H}_{\rm PS}$ $=$ $\frac{({\bm{P}}_{\rm n}-\hat{\bm{p}}_{\rm e})^{2}}{2M}+\hat{H}_{\rm el},$ (22)
Henceforward, when we speak of phase-space electronic Hamiltonians, we will refer to Hamiltonians similar to Eq. 22 above. Note that the eigenvalues of Eq. 22 are identical to the exact energies (up to a uniform constant shift) for a hydrogen atom
$E_{n}^{\rm PS}({\bm{P}}_{\rm n})$ $=$ $\frac{\bm{P}_{\rm n}^{2}}{2(M+m_{\rm e})}-\frac{\mu e^{4}}{8\epsilon_{0}^{2}h^{2}}\frac{1}{n^{2}}$ (23)
provided we equate ${\bm{P}}_{\rm n}$ (which captures the nuclear momentum) with $\bm{P}$ (which is formally the center of mass momentum).
The fact that we can generate the exact eigenvalues through Eq. 22 is very encouraging because the steps taken above offer an approach to generate phase-space Hamiltonians for systems with $N$ isolated hydrogen atoms. In such a case, all we must do is to partition the electronic momentum into different operators, one around each atom, so that one can effectively perform $N$ different local changes of coordinate around each atom $A$:
@@
Here, we define $\hat{\theta}_{A}(\hat{\bm{r}})$ to be a partition of space:
@@
Note that while there are many ways to choose the space partition function $\hat{\theta}_{A}$, a natural option is to choose smooth functions that are localized on atoms (e.g., a combination of Gaussian functions).
One can then make the ansatz that, for a general system with many electrons and nuclei, a phase-space electronic Hamiltonian will be of the form:
@@
To our knowledge, Eqs. 27 and 28 offer the simplest form for a many-electron/many-nuclei phase-space electronic Hamiltonian with clear physical meaning.
2.2 Phase-Space Electronic Hamiltonians and a Simple Form of $\hat{\bm{\Gamma}}_{A}$
Eq. 27 above was guessed on the basis of the hydrogen atom. To better understand where such an equation comes from, and to better improve upon it, note that, if one diagonalizes $\hat{H}_{\rm el}\left(\bm{R}\right)$ in Eq. 27, the exact Hamiltonian in the basis of electronic eigenstates {|$\{\ket{\psi_{j}}\}$ is:
@@ 2.2 Phase-Space Electronic Hamiltonians and a Simple Form of $\hat{\bm{\Gamma}}_{A}$
where $\hat{\bm{d}}_{A}=\sum_{jk}{\bm{d}}_{jk}^{A}\ket{\psi_{j}}\bra{\psi_{k}}$ | is the derivative coupling operator and $\hat{V}_{\rm ad}$ is the diagonal matrix of eigenvalues from Eq. 12. Thus, a phase-space approach is an attempt to capture some of the beyond BO physics present in Eq. 29 by finding an operator $\hat{\bm{\Gamma}}$ that mimics the derivative coupling $\hat{\bm{d}}$.
Now, formally, the derivative coupling operator $\hat{\bm{d}}_{A}$ is not an operator so much as a (expensive) matrix of elements in the adiabatic basis. That being said, it is not difficult to show that ${\bm{d}}^{A}_{jk}$ satisfies four properties:
$-i\hbar\sum_{A}{\bm{d}}^{A}_{jk}+\bra{\psi_{j}}\hat{\bm{p}}\ket{\psi_{k}}$ $=$ $0,$ (30)
$\sum_{A}\bm{\nabla}_{A}{\bm{d}}^{B}_{jk}$ $=$ 00 0 (31)
$-i\hbar\sum_{A}{\bm{R}}_{A}\times{\bm{d}}^{A}_{jk}+\bra{\psi_{j}}\hat{\bm{l}}+\hat{\bm{s}}\ket{\psi_{k}}$ $=$ $0,$ (32)
$-\sum_{A}{\bm{R}}_{A}\times\bm{\nabla}_{A}{d}^{B\beta}_{jk}$ $=$ $\sum_{\gamma}\epsilon_{\alpha\beta\gamma}d^{B\gamma}_{jk},$ (33)
where $\hat{\bm{l}}$ and $\hat{\bm{s}}$ are the electronic orbital and spin angular momentum operators. In words, Eqs. 30 and 32 reflect the fact that the system is unchanged under total translation and rotation of the entire system (of electrons and nuclei). Eqs. 31 and 33 reflect that the derivative coupling vectors themselves transform correctly under translation and rotation.
Now, in developing a meaningful phase-space approach of the form
$\hat{H}_{\rm PS}$ $=$ $\sum_{A}\frac{{\left(\bm{P}_{A}-i\hbar\hat{\bm{\Gamma}}_{A}\right)}^{2}}{2M_{A}}+\hat{H}_{\rm el}({\bm{R}}),$ (34)
one criterion for constructing a meaningful $\hat{\bm{\Gamma}}$ operator is that the operator should satisfy the parallel constraints: